OpenStudy (anonymous):

Find the equation of the line y=-7 and passing through the midpoint of the segment joining (-7, 7) and (-4, 4).

OpenStudy (anonymous):

I'm getting y=-x but webassign says I'm wrong...

OpenStudy (anonymous):

ur question is not clear! what is "find the equation of the line y=-7..."..y=-7 is itself a line equation..

OpenStudy (anonymous):

I know that but this a question from my online math...

OpenStudy (anonymous):

can u gimme a link?

OpenStudy (anonymous):

I don't think so because I have to log into an account

OpenStudy (anonymous):

Find the equation of the line parallel to the line y = -7 and passing through the midpoint of the segment joining (-7, 7) and (-4, 4). I can copy and paste it.

OpenStudy (anonymous):

LOL, I see what I did, I left a few words out, sorry!

OpenStudy (anonymous):

A line parallel to y=-7 is y=c where c is a constant to be determined. NOw midpoint of (-7,7) and (-4,4) is (-11/2,11,2). Now y=c passes through this point. So c=11/2. and answer is y=11/2

OpenStudy (anonymous):

Okay, thanks so much! Can you show me how it's worked out? I want to make sure I understand.

OpenStudy (anonymous):

Ohh okay. NOW your question makes sense but it was already answered : p

OpenStudy (anonymous):

Sorry Mr. H! I forgot a few words when typing!

OpenStudy (anonymous):

where do u get the problem in my solution, bgauldin? :)

OpenStudy (anonymous):

no problem in the solution, I was just wanting to know how to work it out....